OpenStudy (anonymous):

find the simple interest. then find balance principal= $870 interest=3% time=9 moths

OpenStudy (paki):

870*3*9/100 = 235$

OpenStudy (anonymous):

whats the simple interest @paki

OpenStudy (paki):

Simple interest is generally charged for borrowing money for short periods of time. Compound interest is similar but the total amount due at the end of each period is calculated and further interest is charged against both the original principal but also the interest that was earned during that period.... @noah4151

OpenStudy (anonymous):

i knowk but what would it be in the problem @paki

OpenStudy (tkhunny):

Put this in your head... I = P*r*t Fill in the parts you know and find the part that is left. P = principal = $870 r = interest = 3% = 0.03 t = time = 9 moths = 3/4 year I = ($870)(0.03)(3/4) Calculator time!

OpenStudy (tkhunny):

It's way less than $235. @paki neglected to convert the time to years to be consistent with the interest rate.